OpenStudy (kitkat1):

Choose the correct slope of the line that passes through the points (1, −3) and (3, −5). (1 point) −1 1 0 one half

OpenStudy (gabylovesyou):

The equation for slope is y2 - y1 / x2 - x1 Put ur points in the equation to get -5 - (-3) / 3 - 1 -2 / 2 = -1

OpenStudy (netlopes1):

Please look image "kitkat1" I'm finishing ok? Can you wait just a moment?

OpenStudy (netlopes1):

The final resolution is in the attached file, ok?

OpenStudy (kitkat1):

okay thanks can u help me with one more?

OpenStudy (kitkat1):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). (2 points) y − 4 = − one half(x + 6) y − 4 = 2(x + 6) y + 6 = − one half(x − 4) y + 6 = 2(x − 4)

OpenStudy (netlopes1):

ok... m = [0-(-4)] /[2-(-6)] or m=4/8 or m=1/2 ou m=0.5 ... still r: mx+n, where m=0.5

OpenStudy (netlopes1):

Yes, i must find "n", ok?

OpenStudy (netlopes1):

\[I\quad use\quad (-6,4)\quad y=mx+n\rightarrow4=0.5*(-6)+n\rightarrow n=7\]

OpenStudy (netlopes1):

please, i wrong.... m=-0.5

OpenStudy (netlopes1):

The answer is "A" or\[y-4=-0.5(x+6)\]

OpenStudy (netlopes1):

The answer is "A", but how I wrong, if you want, I can (re)make the question and to attach a file (.png), ok?